# RAG-Driver: Generalisable Driving Explanations with Retrieval-Augmented In-Context Learning in Multi-Modal Large Language Model

### v1

@@ III-B Training Strategy
Following the visual instruction tuning paradigm [48, 46], we employ a two-stage training strategy to progressively enable cross-modality alignment and multi-task driving capacity. In both stages, we leverage the same next-token prediction cross-entropy loss as in Eq. 3 aiming to maximize the conditional answer likelihood in Eq. 2, where $\bm{y}$ is the ground truth token.
@@
Pre-training In the first pre-training stage, we only train the cross-modality projector while freezing the visual encoder and LLM. We use a subset VIDAL-10M [85] which contains 3 million video-caption pairs. This achieves an alignment between visual and language features by projecting the pre-trained video embeddings into language tokens understandable by LLM.
Supervised In-context Instruction Tuning While state-of-the-art LLM exhibits zero-shot ICL capability, several works [18] and our ablation study Sec. IV-D have shown further improvement when specifically trained on curated ICL demonstrations. In particular, we construct a multi-modal instruction tuning dataset with structured ICL examples based on the BDD-X dataset [38] resulting in 16K video question and answer pairs. As shown in Fig. 3, for an 8-frame driving video sequence with associated control signals—speed, course, acceleration, and curvature as the current query, we use the retrieval mechanism as in Sec. III-C to retrieve 2 relevant driving experiences, which are then prefixed to the current query as ICL examples. The dataset is tailored to support three distinct tasks, each represented through question-answer pairs in natural language. Note that, with (1) Action Explanation and (2) Justification naturally represented as natural language, (3) Control Signal prediction is also formed as language token prediction; this is feasible due to the distinct mapping of numerical values to specific tokens within the language model dictionary.
@@ III-C Retrival-Augmented In-context Learning
Retrieval Mechanism To perform the retrieval, we first leverage a lightweight MLP projector of the same structure as in Eq. 1 to project the heterogeneous video and control signal embedding into the same hybrid embedding $\bm{s}\in\mathbb{R}^{1024}$ through metric learning [27]. In particular, we adopt a triplet loss with Euclidean distance as shown in Eq. 4:
@@
Retrieval Augmented In-Context Learning (RA-ICL) To perform the RA-ICL, we prefix retrieved samples before the current query, facilitating an implicit gradient descent through the meta-optimiser of LLM, as proved in [17]. This approach is also applicable to MLLM with architecture specified in Sec. III-A. For a given transformer-based pre-trained MLLM modeling the prefix-conditioned output probability as in Eq. 2, consider one head of the multi-head attention in a single transformer block as follows:
@@
where $\bm{z}_{1:n}=\left[\bm{z}_{icl};\bm{z}_{q}\right]$ represented the conditional prefix consist with ICL example $\bm{z}_{icl}$ and current query embeddings $\bm{z}_{q}$, respectively. $W_{Q,K,V}\in\mathbb{R}^{d^{i}\times d^{o}}$ is the linear transformation on query, key, and value in attention block, respectively. Now, in [17] in-context learning was shown to effectively achieve a meta-optimisation to implicitly update the model with an estimated meta-gradient. We provide in our work a new, alternative derivation of this, supported by more recent work in [40] on the following softmax-free linear attention expressions
@@ IV-A Settings and Datasets
We empirically evaluate the proposed Retrieval-augmented In-Context Learning (RA-ICL) framework within the Multi-modal Large Language Model (MLLM), targeting explainable driving applications. We aim to validate its efficacy in general driving scenarios with a focus on two main aspects: (1) explainability in driving action explanation and justification. (2) Control Signal Prediction. We conduct experiments with the BDD-X [38] dataset, which is a widely adopted benchmark in explainable driving, comprising 77-hour videos across the US under different road and weather conditions. We customize the format as shown in Fig. 3, resulting in 16,803 and 2,123 video question-answering pairs for training and testing, respectively. More importantly, we further explore the transfer learning capacity of zero-shot generalisation in unseen environments. We leverage customised dataset comprising 58 testing question-answering pairs, recorded in London, UK, presenting a significant distribution shift from the BDD-X dataset.
Benchmark Settings For all experiments, we train the MLLM using the BDD-X training split. Subsequent evaluation on general explainability and control signal prediction capabilities tests are conducted on the BDD-X test split with the BDD-X training split as the memory database. For the transfer learning experiments, we employ the same foundational model and test it on the Spoken-SAX but the memory database is constructed using the BDD-X training split for zero-shot generalisation.
@@
S2VT [74] $\times$ $30.2$ $179.8$ $27.5$ $6.3$ $53.4$ $11.2$
S2VT++ [74] $\times$ $27.1$ $157.0$ $26.4$ $5.8$ $52.7$ $10.9$
SAA [38] $\times$ $31.8$ $214.8$ $29.1$ 7.1 $66.1$ $12.2$
WAA [38] $\times$ $32.3$ $215.8$ $29.2$ $7.3$ $69.5$ $12.2$
ADAPT [33] $\times$ $34.6$ $247.5$ $30.6$ $11.4$ $102.6$ $15.2$
DriveGPT4 [78] $30.0$ $214.0$ $29.8$ $9.4$ $102.7$ $14.6$
@@
Zero-shot Generalisation (Spoken-SAX)
ADAPT [33] $\times$ $0.0$ $22.3$ $26.7$ $3.1$ $8.6$ $12.4$
@@
ADAPT [33] $5.87$ $54.49$ $86.39$ $91.06$ $97.36$ $98.20$ $2.68$ $11.77$ $31.79$ $47.48$ $92.75$ $95.87$
DriveGPT4 [78] $4.57$ $69.22$ $79.14$ $84.47$ $95.72$ $96.74$ $1.09$ 56.93 $77.77$ $87.97$ $99.00$ $99.57$
@@
Figure 4: Qualitative Examples. Comparison between human annotated ground truth (Left) and prediction from our method (Right) from both BDD-X (Top 4) and Spoken-London (Bottom 4). Each video is downsampled to four frames for visualisation. Despite subtle differences in wording, our method provides driving action explanations and justifications that are comprehensible to humans.
Baselines We compare with various driving action description and justification baseline methods such as video-language sequence-to-sequence based recurrent neural network S2VT [74] and visual-attention based convolutional neural network WAA [38]. Also, we compare with the state-of-the-art explainable driving methods, including video transformer-based ADAPT [33] and visual instruction tuned MLLM DriveGPT4 [78] that is further capable of control signal prediction.
@@ IV-B Explainability in Driving Action and Justification
We begin by evaluating the quality and accuracy of explanations and justifications for driving actions separately. As shown in upper part of Tab. I, our method demonstrates comparable performance to the state-of-the-art specialist method ADAPT [33], a characteristic not observed in previous MLLM-based methods. In particular, when compared with DriveGPT4 [78] which also uses MLLM with a similar architecture and number of parameters but incorporates the extra LLaVA-150K dataset [48] for visual instruction tuning, our approach, relying solely on the BDD-X dataset, outperforms it in terms of explainability. This is evidenced by an average performance improvement of 10.8% across all metrics. This underscores the effectiveness of ICL in enhancing the emergent reasoning capabilities of MLLMs.
@@ IV-D Ablation Study on Retrieval Strategy
We also investigate whether to apply ICL examples during training or solely in inference time. As shown in Tab. III, we found that the MLLM is incapable of making reasonable predictions using ICL examples without prior training, regardless of the retrieval strategy chosen. This suggests that the pre-trained MLLM is not equipped to effectively perform zero-shot ICL. We hypothesize that supervised fine-tuning plays a crucial role in enhancing the ICL capabilities of MLLM, necessitating the provision of reasoning demonstrations, which aligns with observations in [18].
@@ IV-F Qualitative Demonstration
We also demonstrate a series of quantitative examples comparing the driving action explanation and justification provided by human ground truth and prediction from our method. As shown in Fig. 4, we observe RAG-Driver produces robust intelligible action explanation and justification under different environments (i.e. night time and adversarial weather) with a control signal close to the human driver record. More importantly, in the out-of-distribution setting Spoken-SAX as indicated by the clear visual difference, we observe the prediction also produces human-understandable answers, qualitatively validating the exceptional zero-shot generalisation capacity.
@@ V Limitations and Future Work
This work aims to develop a generalisable explainable driving commentator using a Machine Learning Language Model (MLLM), addressing a significant obstacle that has hindered deployment: the poor generalisation capacity. However, several problems still need to be addressed. For instance, the prevalent issue of hallucination in MLLM, while mitigated, is still observed. We hypothesize that this is due to the limited general video understanding capacity, as the video encoder only processes 8 frames per video. Also, due to the limited access to open-source models and computational cost, we employ a relatively small MLLM with 7 billion parameters, which falls short of some state-of-the-art models (i.e. GPT4V [1], Gemini [70]). We anticipate that the development of a more powerful MLLM backbone with lower computational costs could further enhance the application of MLLM in driving.

### v2

@@ III-B Training Strategy
Following the visual instruction tuning paradigm [liu2023llava, lin2023video], we employ a two-stage training strategy to progressively enable cross-modality alignment and multi-task driving capacity. In both stages, we leverage the same next-token prediction cross-entropy loss as in Eq. 3 aiming to maximize the conditional answer likelihood in Eq. 2, where $\bm{y}$ is the ground truth token.
@@
Pre-training In the first pre-training stage, we only train the cross-modality projector while freezing the visual encoder and LLM. We use a subset VIDAL-10M [zhu2023languagebind] which contains 3 million video-caption pairs. This achieves an alignment between visual and language features by projecting the pre-trained video embeddings into language tokens understandable by LLM.
Supervised In-context Instruction Tuning While state-of-the-art LLM exhibits zero-shot ICL capability, several works [dong2022survey] and our ablation study Sec. IV-D have shown further improvement when specifically trained on curated ICL demonstrations. In particular, we construct a multi-modal instruction tuning dataset with structured ICL examples based on the BDD-X dataset [kim2018bddx] resulting in 16K video question and answer pairs. As shown in Fig. 3, for an 8-frame driving video sequence with associated control signals—speed, course, acceleration, and curvature as the current query, we use the retrieval mechanism as in Sec. III-C to retrieve 2 relevant driving experiences, which are then prefixed to the current query as ICL examples. The dataset is tailored to support three distinct tasks, each represented through question-answer pairs in natural language. Note that, with (1) Action Explanation and (2) Justification naturally represented as natural language, (3) Control Signal prediction is also formed as language token prediction; this is feasible due to the distinct mapping of numerical values to specific tokens within the language model dictionary.
@@ III-C Retrival-Augmented In-context Learning
Retrieval Mechanism To perform the retrieval, we first leverage a lightweight MLP projector of the same structure as in Eq. 1 to project the heterogeneous video and control signal embedding into the same hybrid embedding $\bm{s}\in\mathbb{R}^{1024}$ through metric learning [hoffer2015deep]. In particular, we adopt a triplet loss with Euclidean distance as shown in Eq. 4:
@@
Retrieval Augmented In-Context Learning (RA-ICL) To perform the RA-ICL, we prefix retrieved samples before the current query, facilitating an implicit gradient descent through the meta-optimiser of LLM, as proved in [dai2022can]. This approach is also applicable to MLLM with architecture specified in Sec. III-A. For a given transformer-based pre-trained MLLM modeling the prefix-conditioned output probability as in Eq. 2, consider one head of the multi-head attention in a single transformer block as follows:
@@
where $\bm{z}_{1:n}=\left[\bm{z}_{icl};\bm{z}_{q}\right]$ represented the conditional prefix consist with ICL example $\bm{z}_{icl}$ and current query embeddings $\bm{z}_{q}$, respectively. $W_{Q,K,V}\in\mathbb{R}^{d^{i}\times d^{o}}$ is the linear transformation on query, key, and value in attention block, respectively. Now, in [dai2022can] in-context learning was shown to effectively achieve a meta-optimisation to implicitly update the model with an estimated meta-gradient. We provide in our work a new, alternative derivation of this, supported by more recent work in [koohpayegani2024sima] on the following softmax-free linear attention expressions
@@ IV-A Settings and Datasets
We empirically evaluate the proposed Retrieval-augmented In-Context Learning (RA-ICL) framework within the Multi-modal Large Language Model (MLLM), targeting explainable driving applications. We aim to validate its efficacy in general driving scenarios with a focus on two main aspects: (1) explainability in driving action explanation and justification. (2) Control Signal Prediction. We conduct experiments with the BDD-X [kim2018bddx] dataset, which is a widely adopted benchmark in explainable driving, comprising 77-hour videos across the US under different road and weather conditions. We customize the format as shown in Fig. 3, resulting in 16,803 and 2,123 video question-answering pairs for training and testing, respectively. More importantly, we further explore the transfer learning capacity of zero-shot generalisation in unseen environments. We leverage customised dataset comprising 58 testing question-answering pairs, recorded in London, UK, presenting a significant distribution shift from the BDD-X dataset.
Benchmark Settings For all experiments, we train the MLLM using the BDD-X training split. Subsequent evaluation on general explainability and control signal prediction capabilities tests are conducted on the BDD-X test split with the BDD-X training split as the memory database. For the transfer learning experiments, we employ the same foundational model and test it on the Spoken-London but the memory database is constructed using the BDD-X training split for zero-shot generalisation.
@@
S2VT [venugopalan2015sequence] $\times$ $30.2$ $179.8$ $27.5$ $6.3$ $53.4$ $11.2$
S2VT++ [venugopalan2015sequence] $\times$ $27.1$ $157.0$ $26.4$ $5.8$ $52.7$ $10.9$
SAA [kim2018bddx] $\times$ $31.8$ $214.8$ $29.1$ 7.1 $66.1$ $12.2$
WAA [kim2018bddx] $\times$ $32.3$ $215.8$ $29.2$ $7.3$ $69.5$ $12.2$
ADAPT [jin2023adapt] $\times$ $34.6$ $247.5$ $30.6$ $11.4$ $102.6$ $15.2$
DriveGPT4 [xu2023drivegpt4] $30.0$ $214.0$ $29.8$ $9.4$ $102.7$ $14.6$
@@
Zero-shot Generalisation (Spoken-SAX) B4 $\uparrow$ C $\uparrow$ M $\uparrow$ B3 $\uparrow$ C $\uparrow$ M $\uparrow$
ADAPT [jin2023adapt] $\times$ $0.0$ $22.3$ $26.7$ $3.1$ $8.6$ $12.4$
@@
ADAPT [jin2023adapt] $5.87$ $54.49$ $86.39$ $91.06$ $97.36$ $98.20$ $2.68$ $11.77$ $31.79$ $47.48$ $92.75$ $95.87$
DriveGPT4 [xu2023drivegpt4] $4.57$ $69.22$ $79.14$ $84.47$ $95.72$ $96.74$ $1.09$ 56.93 $77.77$ $87.97$ $99.00$ $99.57$
@@
[…]
Baselines We compare with various driving action description and justification baseline methods such as video-language sequence-to-sequence based recurrent neural network S2VT [venugopalan2015sequence] and visual-attention based convolutional neural network WAA [kim2018bddx]. Also, we compare with the state-of-the-art explainable driving methods, including video transformer-based ADAPT [jin2023adapt] and visual instruction tuned MLLM DriveGPT4 [xu2023drivegpt4] that is further capable of control signal prediction.
@@ IV-B Explainability in Driving Action and Justification
We begin by evaluating the quality and accuracy of explanations and justifications for driving actions separately. As shown in upper part of Tab. I, our method demonstrates comparable performance to the state-of-the-art specialist method ADAPT [jin2023adapt], a characteristic not observed in previous MLLM-based methods. In particular, when compared with DriveGPT4 [xu2023drivegpt4] which also uses MLLM with a similar architecture and number of parameters but incorporates the extra LLaVA-150K dataset [liu2023llava] for visual instruction tuning, our approach, relying solely on the BDD-X dataset, outperforms it in terms of explainability. This is evidenced by an average performance improvement of 10.8% across all metrics. This underscores the effectiveness of ICL in enhancing the emergent reasoning capabilities of MLLMs.
@@ IV-D Ablation Study on Retrieval Strategy
We also investigate whether to apply ICL examples during training or solely in inference time. As shown in Tab. III, we found that the MLLM is incapable of making reasonable predictions using ICL examples without prior training, regardless of the retrieval strategy chosen. This suggests that the pre-trained MLLM is not equipped to effectively perform zero-shot ICL. We hypothesize that supervised fine-tuning plays a crucial role in enhancing the ICL capabilities of MLLM, necessitating the provision of reasoning demonstrations, which aligns with observations in [dong2022survey].
@@ IV-F Qualitative Demonstration
We also demonstrate a series of quantitative examples comparing the driving action explanation and justification provided by human ground truth and prediction from our method. As shown in Fig. 4, we observe RAG-Driver produces robust intelligible action explanation and justification under different environments (i.e. night time and adversarial weather) with a control signal close to the human driver record. More importantly, in the out-of-distribution setting Spoken-London as indicated by the clear visual difference, we observe the prediction also produces human-understandable answers, qualitatively validating the exceptional zero-shot generalisation capacity.
@@ V Limitations and Future Work
This work aims to develop a generalisable explainable driving commentator using a Machine Learning Language Model (MLLM), addressing a significant obstacle that has hindered deployment: the poor generalisation capacity. However, several problems still need to be addressed. For instance, the prevalent issue of hallucination in MLLM, while mitigated, is still observed. We hypothesize that this is due to the limited general video understanding capacity, as the video encoder only processes 8 frames per video. Also, due to the limited access to open-source models and computational cost, we employ a relatively small MLLM with 7 billion parameters, which falls short of some state-of-the-art models (i.e. GPT4V [achiam2023gpt], Gemini [team2023gemini]). We anticipate that the development of a more powerful MLLM backbone with lower computational costs could further enhance the application of MLLM in driving.
@@ VI Conclusion
-A Discussion and Limitation
Scale of MLLM While our Multi-modal Large Language Model (MLLM) has exhibited impressive capabilities in visual reasoning and planning for driving tasks, it’s worth noting that it comprises only 7 billion parameters. This size is relatively modest when compared to more well-known models such as GPT4-V [achiam2023gpt] and Gemini [team2023gemini], which boast a significantly larger parameter count and demonstrate superior, near-human levels of visual understanding and reasoning. In various related fields such as visual question answering, problem-solving, and interactive dialogue, a clear trend has been observed by researchers: the scale of the model’s parameters and the breadth of training data sources are pivotal. Performance improvements are typically seen in tandem with model scaling. Based on this trend, we expect similar advancements in the realm of driving applications. A larger model could further enhance the capabilities of MLLM in driving scenarios.
Number of In-Context Learning Examples During training and inference, we provide 2 ICL examples for each query. This is due to the limitation of the context window size of the LLM backbone of 2048. With the recent improvement in LLM context window size, we expect to see a more flexible adoption of ICL examples.
-B Training Details
Embedding Projector We leverage a three-layer MLP as the embedding projector to fuse the video $1\times 1024$ and control signal $1\times 28$ embedding into a hybrid embedding. The architecture of the lightweight projector is a four-layer MLP with GELU activation. It has an input dimension of $1052$ and an output dimension of $1024$. The margin used in triplet loss is 0.5. We train the model for 200 epochs with a learning rate of $51e-$ 5 with Adam optimiser.
MLLM Backbone We use a learning rate of $52e-$ 5 with a cosine scheduler. We use a batch size of 4 with a gradient accumulation step of 2 on 8 A100 GPUs, which leads to an effective training batch size of 128. We use a warm-up strategy in the first 5 epochs with a warm-up ratio of 0.03. We train the model for 2 epochs.
-C Baseline Details
In our comparison, we evaluate several baseline methods. The first, S2VT [venugopalan2015sequence], utilises an end-to-end sequence-to-sequence model with Long Short-Term Memory (LSTM) networks. It is trained on paired video-sentence data, linking video frame sequences to corresponding word sequences, enabling it to generate descriptive captions for video events. The second method, WAA [kim2018bddx], employs a visual attention model that trains a convolutional network from images to vehicle control commands. This method focuses on identifying influential image regions through the controller’s attention and uses an attention-based video-to-text model to produce textual explanations aligned with the controller’s attention maps, grounding the explanations in relevant scene parts. The third approach, ADAPT [jin2023adapt], is a transformer-based method that leverages a multi-task joint training framework. It aligns driving action captioning with control signal prediction tasks. Finally, DriveGPT4 [xu2023drivegpt4], trained using LLaVA [liu2023llava] to generate a visual instruction tuning dataset derived from the BDD-X dataset [kim2018bddx], processes multimodal input data and is capable of generating text responses while predicting control signals, fine-tuned with the assistance of ChatGPT on the new dataset.
-D Linear Layer Parameter Update Derivations
Consider a forward-pass through a linear layer $\mathcal{F}$ after its weights $W_{0}$ have been updated by $\Delta{}W$
$\mathbf{y}=\mathcal{F}(\mathbf{x})=(W_{0}+\Delta{}W)\mathbf{x}=W_{0}\mathbf{x}+\Delta{}W\mathbf{x}$
The weight update itself is expressed as
$\Delta{}W=\eta\frac{\partial{}L}{\partial\mathbf{y}}|_{\mathbf{y}_{i}}\mathbf{x}_{i}^{T}\Rightarrow\Delta{}W\mathbf{x}=\eta\frac{\partial{}L}{\partial\mathbf{y}}|_{\mathbf{y}_{i}}\mathbf{x}_{i}^{T}\mathbf{x}$
where $\mathbf{x}_{i},\mathbf{y}_{i}$ are the input and output to the layer that resulted in the weight update. Now, if we in fact have optimised over a mini-batch of input-outputs $\mathbf{x}_{i},\mathbf{y}_{i}$ we have
$\Delta{}W=\sum_{i}\eta\frac{\partial{}L}{\partial\mathbf{y}}|_{\mathbf{y}_{i}}\mathbf{x}_{i}^{T}\Rightarrow\Delta{}W\mathbf{x}=\sum_{i}\eta\frac{\partial{}L}{\partial\mathbf{y}}|_{\mathbf{y}_{i}}\mathbf{x}_{i}^{T}\mathbf{x}$
Therefore we have a weighted sum of dot products, which is akin to an attention mechanism. Indeed, from Eq. 6 we can apply softmax-free linear attention expressions such as in [koohpayegani2024sima] for
$W_{V}\left[\bm{z}_{icl};\bm{z}_{q}\right]\mathrm{softmax}\left(\frac{\left(W_{K}\bm{z}_{1:n}\right)^{T}(W_{Q}\bm{z}_{1:n})}{\sqrt{d^{i}}}\right)$
$\rightarrow W_{V}\left[\bm{z}_{icl};\bm{z}_{q}\right]\left(W_{K}\left[\bm{z}_{icl};\bm{z}_{q}\right]\right)^{T}W_{Q}\bm{z}_{1:n}$
Multiplying the linear attention matrices through the stacked in-context and query embeddings we have
$W_{V}\bm{z}_{icl}(W_{K}\bm{z}_{icl})^{T}W_{Q}\bm{z}_{1:n}+W_{V}\bm{z}_{q}(W_{K}\bm{z}_{q})^{T}W_{Q}\bm{z}_{1:n}$
Now take out a common factor of $W_{Q}\bm{z}_{1:n}$ for
$(W_{V}\bm{z}_{icl}(W_{K}\bm{z}_{icl})^{T}+W_{V}\bm{z}_{q}(W_{K}\bm{z}_{q})^{T})W_{Q}\bm{z}_{1:n}$
and put $\Delta{}W_{ICL}=W_{V}\bm{z}_{icl}(W_{K}\bm{z}_{icl})^{T}$ and $W_{ZSL}=W_{V}\bm{z}_{q}(W_{K}\bm{z}_{q})^{T}$ as the terms both pre-multiplying $W_{Q}\bm{z}_{1:n}$. Note that $W_{ZSL}$ is independent of the in-context terms (depending only on the query). Now, in $\Delta{}W_{ICL}$ we in fact have a set of in-context retrieved samples $\bm{z}_{icl,i}$ such that
$W_{V}\left[\bm{z}_{icl,0},\bm{z}_{icl,1},\ldots\right](W_{K}\left[\bm{z}_{icl,0},\bm{z}_{icl,1},\ldots\right])^{T}$
$\rightarrow\Delta{}W_{ICL}=\sum_{i}~{}W_{V}\bm{z}_{icl,i}(W_{K}\bm{z}_{icl,i})^{T}$
Finally, by inspection of similar dot-product expressions $W_{V}\bm{z}_{icl,i}(W_{K}\bm{z}_{icl,i})^{T}\leftrightarrow\eta\frac{\partial{}L}{\partial\mathbf{y}}|_{\mathbf{y}_{i}}\mathbf{x}_{i}^{T}\mathbf{x}$ we note that we match the form for the linear layer above
This can therefore be interpreted to say that the output of the attention is adjusted in a meta-optimal way to conform to the samples provided as input context, much like gradient descent on the linear layer would adjust that layer to conform to the mini-batch training data, but crucially in the case of RAG-Driver, without supervision.